OpenStudy (anonymous):

The volume of a child's model plane is 1200 cm^3 The volume of the full size plane is 4050 m^3 Find the scale of the model in the form 1:n

jhonyy9 (jhonyy9):

will be 4050/1200 = 405/120 = 81/24 = 27/8 = 1/n 27/8 = 1/n cross multiplie 27n=8 n= 8/27 result this scale will be 1/(8/27) hope is understandably

OpenStudy (anonymous):

but 4050 m3 should be changed to cm3 right?

OpenStudy (anonymous):

@jhonyy9

jhonyy9 (jhonyy9):

oh yes sorry i missed it

jhonyy9 (jhonyy9):

or invers 1200 cm^3 = ? m^3

jhonyy9 (jhonyy9):

1 m^3 = ? cm^3

OpenStudy (anonymous):

=1,000,000 cm^3

OpenStudy (anonymous):

so 4050m^3 would be 4,050,000,000?

jhonyy9 (jhonyy9):

yes right but i think will be more easy without this so great numbers than you change 1200 cm^3 in m^3 how you think it ?

jhonyy9 (jhonyy9):

but you can rewrite this result using 4*10^9 (standard form)

jhonyy9 (jhonyy9):

so 4050m^3 would be 4,050,000,000

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

so 4050m^3 would be 4,050,000,000 = 4*10^9 yes ?

jhonyy9 (jhonyy9):

or 1200 cm^3 = ? m^3

OpenStudy (anonymous):

1200 cm^3=0.0012 m^3

jhonyy9 (jhonyy9):

rewrite it using power of 10

jhonyy9 (jhonyy9):

1200 = aproximatly 1*10^3

OpenStudy (anonymous):

1.2x10^-3

OpenStudy (anonymous):

n=4*10^9 / 10^3

jhonyy9 (jhonyy9):

yes 1,2*10^-3 but you can consider it 1*10^-3 = 10^-3 ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

so than will result 1/n = 10^-3 / 4*10^9 n=?

jhonyy9 (jhonyy9):

so n= ?

OpenStudy (anonymous):

2.5*10^-13?

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

but the answer is supposed to be n=150 :/

jhonyy9 (jhonyy9):

yes but you need writing your answer in this 1/n form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

isnt it supposed to be related to that?

jhonyy9 (jhonyy9):

so than us this 2,5 *10^-13 so than from 2,5/4 you will get 0,6 because 1/150 = 0,006 = 6*10^-3